OpenStudy (anonymous):

A right triangle has a hypotenuse of 30 inches and a first leg of 18 inches. What is the length of the second leg?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

30^2 - 18^2 = b^2 900 - 324 576 = b^2 √576 b = 24

OpenStudy (anonymous):

x^2 = (30^2 -18^2) = 48*12 x = sqrt(48*12) = sqrt(16*4*9) = 4*2*3 = 24 inch